# Gauge-Invariant Longitudinal Modes in the Herwig 7 Electroweak Parton Shower

### v1

@@ 4.1 Single-Resummed Evolution
Figures 4-6 show that, for boson-line branchings, the GI prescription yields a systematically larger single-resummed splitting-tagged rate than the Dawson construction, with a mild enhancement in both the parent-$p_{\perp}$ spectra and the reconstructed energy-sharing variable $z$ (typically at the $\mathcal{O}(5\text{-}15\%)$ level, with the largest deviations at low $p_{\perp}$ and away from the kinematic end-points). The effect is qualitatively similar across the channels $^{\pm}\!\to W^{\pm}Z^{0}, W±W$$^{\pm}\!\to W^{\pm}\gamma$, $Z^{0}\!\to W^{+}W^{-}$ and $\gamma\!\to W^{+}W^{-}$, consistent with the fact that the comparison is dominated by the longitudinally sensitive part of the bosonic $V\!\to V^{\prime}V^{\prime\prime}$ kernels rather than by the hard-process normalisation. The sign relative to the $q\!\to qV$ case is not paradoxical: in the fermion-line study, the analysis selects an exclusive one-boson ensemble, so a locally larger GI rate increases the Sudakov suppression and migrates probability into vetoed higher-multiplicity configurations, which can make the observed GI spectrum smaller even when the underlying kernel is larger. For the boson-line channels in Figs. 4-6, the tagged $V\!\to V^{\prime}V^{\prime\prime}$ branchings contribute less to the Sudakov exponent controlling the acceptance, so the change in overall no-emission probability is comparatively small and the enhanced GI longitudinal contribution (fixed by the Ward-identity/Goldstone completion) manifests directly as an upward shift in the splitting-tagged distributions, hence ${\rm GI}/{\rm Dawson}>1$ over most of the measured phase space.
@@ 4.2 Full-Resummed Evolution
The resulting samples are analysed with dedicated Rivet routines which identify $W/Z$ occurrences in the event record and construct inclusive spectra and reconstructed shower variables for representative bosonic channels. In the full-resummed context, this provides an internal consistency check of the longitudinal completion: observables dominated by last-copy, on-shell gauge bosons mainly probe the choice of physical polarisation basis and are therefore expected to agree between Dawson and GI up to subleading recoil and bookkeeping effects. By contrast, splitting-tagged configurations probe iterated, generally off-shell gauge-boson currents inside the shower, whose longitudinal component is fixed by the broken-theory Ward identity and its associated Goldstone-matching; it is therefore in such samples, particularly in bosonic channels and ultracollinear regions, that the largest sensitivity to the GI completion is expected.
@@ 5 Conclusions
We have constructed and implemented a GI longitudinal completion for EW parton-shower branchings in the broken theory, formulated at the level of helicity-dependent quasi-collinear splitting operators. The construction enforces the Ward-identity constraints on longitudinal currents by explicit Goldstone-matching, thereby removing the scheme ambiguity associated with the naive $p^{\mu}/m_{V}$ growth of $\varepsilon^{\mu}_{L}$ while retaining a probabilistic shower interpretation and spin correlations. We provided the reduced helicity amplitudes for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ in the Herwig 7 conventions and exposed the minimal modification pattern: only the longitudinal $0_{*}$ component receives GI contributions, controlled by the Goldstone coefficient and Yukawa structures, leaving the transverse sector unchanged. In practical terms, this means inclusive observables dominated by promptly contracted, near on-shell vector-boson matrix elements are largely unaffected, while splitting-tagged and reconstructed shower-current observables provide the cleanest handles on the longitudinal scheme. The corresponding implementation is numerically stable and configurable through run-card switches, enabling systematic comparisons with the widely used Dawson prescription and providing a convenient starting point for extensions to non-standard chiral and triple-gauge couplings.
@@
We emphasise that the Dawson-GI difference is parametrically controlled by $m_{1,2}^{2}/\tilde{q}^{\,2}$, so the two prescriptions must coincide at high evolution scales while deviations are expected precisely in the low-$\tilde{q}$ region where symmetry-breaking terms become numerically relevant. In that region, the shower cutoff/physicality constraints and the exclusivity of the one-emission selection can amplify small kernel-level shifts into non-monotonic changes in accepted event samples through Sudakov suppression and vetoed higher-multiplicity configurations.
@@ Acknowledgements
We thank our fellow Herwig 7 authors for useful discussions. This work has received funding from the European Union’s Horizon 2020 research and innovation program as part of the Marie Skłodowska-Curie Innovative Training Network MCnetITN3 (grant agreement no. 722104). MRM is also supported by the UK Science and Technology Facilities Council (grant numbers ST/T001011/1 and ST/X000745/1).
@@ Appendix A Run-card interface for the Longitudinal EW Schemes
with relevant classes being QtoQWZSudakov for $q\to q^{\prime}V$ splittings, VtoVVSudakov for $V\to V^{\prime}V^{\prime\prime}$ splittings, and GammatoWWSudakov for $\gamma\to W^{+}W^{-}$ splitting. The <picture> can be set to either Dawson or GaugeInvariant. In the GI case, the relative coefficient $c_{G}$ multiplying the Goldstone-matching contribution can be configured via
@@
with the default set to $c_{G}=1$. Setting $c_{G}=0$ switches off the Goldstone-matching term, leaving only the Dawson remainder in the longitudinal current. Other values ($0<c_{G}<1$, $c_{G}>1$, or $c_{G}<0$) may be used as a purely diagnostic interpolation/stress test, but they do not correspond to a gauge-invariant SM prescription. In practice, the spread between $c_{G}=1$ and $c_{G}=0$ can be taken as an estimate of the residual scheme/model dependence associated with the treatment of longitudinal emissions: this difference is controlled by symmetry-breaking mass effects and therefore decreases with increasing evolution scale.

### v2

@@ 4.1 Single-Resummed Evolution
Figures 4-6 show that, for boson-line branchings, the GI prescription yields a systematically larger single-resummed splitting-tagged rate than the SL construction, with a mild enhancement in both the parent-$p_{\perp}$ spectra and the reconstructed energy-sharing variable $z$ (typically at the $\mathcal{O}(5\text{-}15\%)$ level, with the largest deviations at low $p_{\perp}$ and away from the kinematic end-points). The effect is qualitatively similar across the channels $^{\pm}\!\to W^{\pm}Z^{0}, W±W$$^{\pm}\!\to W^{\pm}\gamma$, $Z^{0}\!\to W^{+}W^{-}$ and $\gamma\!\to W^{+}W^{-}$, consistent with the fact that the comparison is dominated by the longitudinally sensitive part of the bosonic $V\!\to V^{\prime}V^{\prime\prime}$ kernels rather than by the hard-process normalisation. The sign relative to the $q\!\to qV$ case is not paradoxical: in the fermion-line study, the analysis selects an exclusive one-boson ensemble, so a locally larger GI rate increases the Sudakov suppression and migrates probability into vetoed higher-multiplicity configurations, which can make the observed GI spectrum smaller even when the underlying kernel is larger. For the boson-line channels in Figs. 4-6, the tagged $V\!\to V^{\prime}V^{\prime\prime}$ branchings contribute less to the Sudakov exponent controlling the acceptance, so the change in overall no-emission probability is comparatively small and the enhanced GI longitudinal contribution (fixed by the Ward-identity/Goldstone completion) manifests directly as an upward shift in the splitting-tagged distributions, hence ${\rm GI}/{\rm SL}>1$ over most of the measured phase space.
@@ 4.2 Full-Resummed Evolution
The resulting samples are analysed with dedicated Rivet routines which identify $W/Z$ occurrences in the event record and construct inclusive spectra and reconstructed shower variables (including the energy-sharing variable $z$) for representative bosonic channels. In the full-resummed context, this provides an internal consistency check of the longitudinal completion: observables dominated by last-copy, on-shell gauge bosons mainly probe the choice of physical polarisation basis and are therefore expected to agree between SL and GI up to subleading recoil and bookkeeping effects. By contrast, splitting-tagged configurations probe iterated, generally off-shell gauge-boson currents inside the shower, whose longitudinal component is fixed by the broken-theory Ward identity and its associated Goldstone-matching; it is therefore in such samples, particularly in bosonic channels and ultracollinear regions, that the largest sensitivity to the GI completion is expected.
@@ 5 Conclusions
We have constructed and implemented a GI completion for EW parton-shower branchings in the broken theory, formulated at the level of helicity-dependent quasi-collinear splitting operators. The construction enforces the Ward-identity constraints on longitudinal currents by explicit Goldstone-matching, thereby removing the scheme ambiguity associated with the naive $p^{\mu}/m_{V}$ growth of $\varepsilon^{\mu}_{L}$ while retaining a probabilistic shower interpretation and spin correlations. We provided the reduced helicity amplitudes for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ in the Herwig 7 conventions and exposed the minimal modification pattern: only the longitudinal $0_{*}$ component receives GI contributions, controlled by the Goldstone coefficient and Yukawa structures, leaving the transverse sector unchanged. In practical terms, this means inclusive observables dominated by promptly contracted, near on-shell vector-boson matrix elements are largely unaffected, while splitting-tagged and reconstructed shower-current observables provide the cleanest handles on the scheme choice. The corresponding implementation is numerically stable and configurable through run-card switches, enabling systematic comparisons with the widely used SL prescription and providing a convenient starting point for extensions to non-standard chiral and triple-gauge couplings.
@@
We emphasise that the SL-GI difference is parametrically controlled by $m_{1,2}^{2}/\tilde{q}^{\,2}$, so the two prescriptions must coincide at high evolution scales while deviations are expected in the low-$\tilde{q}$ region where symmetry-breaking terms become numerically relevant. In that region, the shower cutoff/physicality constraints and the exclusivity of the one-emission selection can amplify small kernel-level shifts into non-monotonic changes in accepted event samples through Sudakov suppression and vetoed higher-multiplicity configurations.
@@ Acknowledgements
We thank our fellow Herwig 7 authors for useful discussions. This work has received funding from the European Union’s Horizon 2020 research and innovation programme as part of the Marie Skłodowska-Curie Innovative Training Network MCnetITN3 (grant agreement no. 722104). MRM is also supported by the UK Science and Technology Facilities Council (grant numbers ST/T001011/1 and ST/X000745/1).
@@ Appendix A Run-card interface for the Longitudinal EW Schemes
with relevant classes being QtoQWZSudakov for $q\to q^{\prime}V$ splittings, VtoVVSudakov for $V\to V^{\prime}V^{\prime\prime}$ splittings, and GammatoWWSudakov for $\gamma\to W^{+}W^{-}$ splitting. The <picture> can be set to either Subtraction or GaugeInvariant. In the GI case, the relative coefficient $c_{G}$ multiplying the Goldstone-matching contribution can be configured via
@@
with the default set to $c_{G}=1$. Setting $c_{G}=0$ switches off the Goldstone-matching term, leaving only the subtraction remainder in the longitudinal current. Other values ($0<c_{G}<1$, $c_{G}>1$, or $c_{G}<0$) may be used as a purely diagnostic interpolation/stress test, but they do not correspond to a gauge-invariant SM prescription. In practice, the spread between $c_{G}=1$ and $c_{G}=0$ can be taken as an estimate of the residual scheme/model dependence associated with the treatment of longitudinal emissions: this difference is controlled by symmetry-breaking mass effects and therefore decreases with increasing evolution scale.